THE PERMANENT ELECTRIC DIPOLE MOMENTS FOR THE A^{2}Π AND B^{2}Σ^{+} STATES AND THE HYPERFINE INTERACTIONS IN THE A^{2}Π STATE OF LANTHANUM MONOXIDE

\underline{WILTON L. VIRGO} and TIMOTHY C. STEIMLE, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ, 85287-1604.

The optical Stark effect in the (0,0) A^{2}Π_{3/2} − X^{2}Σ^{+}, A^{2}Π_{1/2} − X^{2}Σ^{+} and B^{2}Σ^{+} − X^{2}Σ^{+} band systems of a lanthanum monoxide, LaO, supersonic molecular beam sample have been analyzed to produce permanent electric dipole moments, μ, for the A^{2}Π_{3/2}, A^{2}Π_{1/2} and B^{2}Σ^{+} states of 1.89(6)D, 2.44(2)D and 0.2(1)D, respectively. The ground state dipole moment was determined some time ago^{1} to be 3.207(11)D. Fine structure splitting in the field free A^{2}Π_{3/2} − X^{2}Σ^{+} and A^{2}Π_{1/2} − X^{2}Σ^{+} spectra was analyzed to produce the magnetic hyperfine spectroscopic parameters a = 233(4) MHz, c = -261(12) MHz, and d = 410(4) MHz. The Group IIIA monoxides, ScO, YO and LaO are prototypical examples of transition metal-oxide bonded molecules. This work on LaO has resulted in the first complete set of dipole moment measurements for the X^{2}Σ^{+}, A^{2}Π_{1/2}, A^{2}Π_{3/2} and B^{2}Σ^{+} states for any of these monoxides. The permanent electric dipole moments for the low-lying states of LaO are exceedingly small, given that they are conventionally viewed as having a La^{+2}O^{−2} charge distribution and exhibiting a large variation amongst the four states.

1. R.F. Seunram, F.J. Lovas, G.T. Fraser and K. Matsumura, J. Chem. Phys. 92, 4724 (1990).